OpenStudy (anonymous):

Solve the system of equations. y = -0.04x2 + 8.3x y = 1.12x + 1.6 A. no solution B. (0.22, 1.85) C. (179.28, 202.39) D. (0.22, 1.85) and (179.28, 202.39)

OpenStudy (anonymous):

y=-0.08+8.3 x Subtr. y= 1.6 +1.12x 0=-1.68+7.18x 7.18x=1.68 x=1.68/7.18=0.23 y=1.6+(1.12*0.23)=1.85 x=0.23 and y=1.85

OpenStudy (kropot72):

The correct solution is D.

OpenStudy (kropot72):

The first equation in the question is this?\[y=-0.04x ^{2}+8.3x\]